you have 80 linear feet of fencing with which to enclose a circular space for a garden find the largest area that can be enclosed with this much fencing and the diameter of the corresponding garden


1
Expert's answer
2022-03-15T10:17:57-0400

C=2πr


r=C/(2π)=80/(6.28)=12.


S=πr2=3.14x12.742=509.6 linear square feet